# Multidimensionality of the Hubble tension: the roles of $\Omega_{m}$ and $\omega_{c}$

### v1

@@ III Datasets and methodology
We consider various combinations of the above datasets, following a rationale that will be discussed later.
@@
Assuming $\Lambda$CDM as the underlying cosmological model, we explicitly investigate on real data the connection between $h$ and $\omega_{c}$ discussed earlier, see Eq. (2). We do so by first setting calibrating $\Omega_{m}$ by using BAO and/or SNeIa measurements, as discussed in Sec. II. Once this is achieved, we calibrate the resulting distance ladder on opposite ends, resulting in $H_{0}$ being pushed to either end of the “Hubble tension range”, $0.67\lesssim h\lesssim 0.74$. On the one hand, we can use the BBN prior on $\omega_{b}$ to infer $r_{d}$ (assuming $\Lambda$CDM), in turn calibrating BAO measurements from the early Universe side (this calibration is eventually transferred to SNeIa measurements in the same range, if they are included). The value of $H_{0}$ inferred from this inverse distance ladder is expected to fall on the low side of the Hubble tension range, i.e. closer to the Planck value Schöneberg et al. (2019, 2022b). On the other hand, the SH0ES prior on $H_{0}$ can be used to directly calibrate either or both BAO and SNeIa measurements from the local Universe side, naturally resulting in a value of $H_{0}$ falling on the high side of the Hubble tension range, i.e. closer to the SH0ES value.
@@
Finally, we further quantify the relation between $\omega_{c}$, $S_{8}$, and $H_{0}$ [see Eqs. (4,14)] by fitting linear relations to the relative variations of both $\omega_{c}$ and $S_{8}$ ($\delta\omega_{c}/\omega_{c}$ and $\delta S_{8}/S_{8}$) as a function of the relative variation of $H_{0}$, $\delta h/h$:
@@ IV Results
We now present the results obtained using the methods and datasets discussed in Sec. III. Constraints on cosmological parameters of interest, including the coefficients $\alpha$ and $\beta$ presented in Eq. (15), are shown in Tab. 1. The following four subsections are each devoted to the comparison between the results obtained adopting two specific combinations of likelihoods discussed previously, each of which will make the mutual correlations between $\omega_{c}$, $S_{8}$, and $H_{0}$ more or less clear, while allowing us to underscore the importance of a reliable $\Omega_{m}$ calibration.
@@ IV.2 BBN+PP+BAO vs BBN+PP+SH0ES
In this case, we are no longer adopting the same dataset to calibrate $\Omega_{m}$ (which earlier was provided in both cases by the BAO+PP combination), which is calibrated by BAO+PP on one side, and by PP on the other. The resulting calibration of $\Omega_{m}$ is broadly consistent, but displays some differences. In particular, from the BBN+PP+BAO combination we find $\Omega_{m}=0.314\pm 0.013$, as previously discussed in Sec. IV.1, whereas from the BBN+PP+SH0ES combination we find the $\approx 0.8\sigma$ larger value of $\Omega_{m}=0.331\pm 0.018$, as can be clearly appreciated in Fig. 2. This is not unexpected, as the fact that the PantheonPlus SNeIa sample appears to prefer slightly larger values of $\Omega_{m}\sim 0.33$ compared to its predecessors Pantheon and JLA is well known and documented in the literature.
@@
The above increase in $H_{0}$ leads to substantial increases in both $\omega_{c}$ and $S_{8}$. In fact, we find that $\omega_{c}$ increases by $1.3\sigma$ from $0.134\pm 0.013$ to $0.159\pm 0.015$: we remark that the presence of the BBN prior in the BBN+PP+SH0ES dataset combination plays a crucial role in determining the latter value, given that PP+SH0ES on their own would only be sensitive to the sum of $\omega_{b}$ and $\omega_{c}$ (more precisely, PP is sensitive to $\Omega_{m}$, so adding the SH0ES information on $H_{0}$ naturally determines $\omega_{m}$), whereas the BBN prior allows to disentangle the baryonic contribution from the DM one. Similarly to $\omega_{c}$, we observe that $S_{8}$ increases from $0.846\pm 0.071$ to $0.980\pm 0.084$. In the case of $S_{8}$, although the central value clearly increases to extremely high values, it would be misleading to quantify the significant of the resulting tension with either CMB or weak lensing observations. In fact, we note that the uncertainty on $S_{8}$ is of order $0.07$-$0.08$, i.e. a factor of $\approx 6$-$7$ larger than that of $S_{8}=0.832\pm 0.013$ as determined by Planck, which would result in a tension formally of low significance ($\lesssim 2\sigma$) in spite of the very high central value. The reason for these large uncertainties is to be sought in the fact that our analysis includes neither CMB nor weak lensing data, both of which are crucial to constrain $S_{8}$. Nevertheless, even in this way we are able to observe a significant upwards shift in $S_{8}$, in line with the analytical expectations laid out previously. Finally, the best-fit values for $\alpha$ and $\beta$ we determine for the BBN+PP+SH0ES combination are $\alpha=2.27$ and $\beta=1.73$.
@@ IV.3 BBN+BAO vs BBN+PP+SH0ES
Driven by the previous results, we now choose to compare two dataset combinations which exacerbate the tension between the respective determinations of $H_{0}$, in order to observe a larger shift in $\omega_{c}$ and $S_{8}$. Specifically, we compare the BBN+BAO versus BBN+PP+SH0ES dataset combinations: compared to the earlier discussion in Sec. IV.3, we have removed the PP dataset from the first combination. A visual summary of our results is given in the corner plot of Fig. 3.
In this case, analogously to Sec. IV.3, we are no longer adopting the same dataset to calibrate $\Omega_{m}$, which is calibrated by the BAO dataset on one side, and the PP dataset on the other. In terms of the resulting value of $\Omega_{m}$, these two datasets display the largest difference among all the ones we discussed. In particular, from the BBN+BAO combination we find $\Omega_{m}=0.292\pm 0.019$, whereas from the BBN+PP+SH0ES combination we find the $\approx 1.5\sigma$ larger value of $\Omega_{m}=0.331\pm 0.018$ discussed previously in Sec. IV.3.
@@ IV.4 BBN+BAO vs BBN+P+SH0ES
In light of all the previous considerations, the final scenario we study is one where $\Omega_{m}$ is still calibrated by two different probes, but in a way which is more consistent than the ones considered in Sec. IV.2 and Sec. IV.3. In both cases, we saw that the use of the PantheonPlus SNeIa catalog was responsible for the higher value of $\Omega_{m}\sim 0.33$ inferred in the respective dataset combination. However, it is known that preceding Pantheon SNeIa catalog preferred lower values of $\Omega_{m}$ compared to PantheonPlus. For this reason, we now choose to compare the BBN+BAO versus BBN+P+SH0ES dataset combinations: compared to the earlier discussion in Sec. IV.3, we have therefore replaced the PP dataset in the second combination with the P one. A visual summary of our results is given in the corner plot of Fig. 4. In some sense, this comparison is the best compromise among all the ones we have considered so far. In fact, this comparison allows for a highly consistent calibration of $\Omega_{m}$, while allowing us to explore the opposite ends of the Hubble tension range precisely because the calibration of $\Omega_{m}$ is achieved via different datasets (BAO and SNeIa), which nonetheless are consistent with each other as far as $\Omega_{m}$ is concerned.
@@ V Discussion
The results presented previously fully confirm on real data the expected correlations between $H_{0}$, $\omega_{c}$, and $S_{8}$ which, we recall increase hand-in-hand. While we have specifically assumed $\Lambda$CDM at early times, the increase in $\omega_{c}$ and $S_{8}$ as $H_{0}$ increases does not depend on the specific model assumed for the early Universe, since it is solely a consequence of constraints on $\Omega_{m}$ imposed from the late-time expansion history. In fact, repeating our analysis in a early Universe-agnostic manner, treating $r_{d}$ as a free parameter and eventually imposing some prior thereon to calibrate the BAO measurements from the early Universe side, would lead to essentially the same conclusions. Therefore, we can expect that any early-time modification to $\Lambda$CDM aiming to solve the Hubble tension will necessarily have to be accompanied by an increase in $\omega_{c}$ and, if the primordial power spectrum is $\Lambda$CDM-like, in $S_{8}$ as well. We also note that we have explicitly assumed that $\Lambda$CDM holds at late times, as is usually done when studying early-time models of new physics – we return to this point the end of the Section.
@@
A very important point which is underscored by our analysis is the key role played by a (consistent or not) $\Omega_{m}$ calibration. As highlighted especially in Sec. IV.2 and Sec. IV.3, an inconsistent calibration of $\Omega_{m}$ enhances the shifts in $\omega_{c}$ discussed in this work, which in turn exacerbates the $S_{8}$ discrepancy. This issue is actually highly relevant at present, given the mild disagreement between the values of $\Omega_{m}$ inferred from SNeIa catalogs: as we have discussed earlier, the Pantheon SNeIa dataset indicated $\Omega_{m}\sim 0.3$, which increased to $\Omega_{m}\sim 0.33$ in the PantheonPlus sample. The more recent DES-Y5 and Union3 SNeIa samples instead indicate $\Omega_{m}\sim 0.35$ Abbott et al. (2024) and $\Omega_{m}\sim 0.36$ Rubin et al. (2023) respectively, see also Ref. Colgáin et al. (2024a) (note that these figures are obtained within $\Lambda$CDM, and potentially increase to much larger values when allowing for an evolving dark energy component): we therefore expect that adopting these SNeIa samples would exacerbate the $\omega_{c}$ and $S_{8}$ shifts we have observed. Finally, while we defer a more detailed exploration of this point elsewhere, we note that these shifts in the values of $\Omega_{m}$ from different samples may be connected to recent discussions on redshift-evolution of cosmological parameters as different redshift are probed (see e.g. Refs. Krishnan et al. (2020, 2021a); Dainotti et al. (2021, 2022a); Colgáin et al. (2022); Dainotti et al. (2022b); Colgáin et al. (2024b, 2023); Jia et al. (2023); Malekjani et al. (2024); Adil et al. (2023); Akarsu et al. (2024b); Colgáin et al. (2024c, d); Jia et al. (2024) for more details).
@@
The above discussions reinforce the urgent need, in the cosmology community, for a calibration of $\Omega_{m}$ which is as reliable and model-independent as possible. Concerning this last point, we indeed note that the value of $\Omega_{m}$ inferred from standard late-time cosmological probes (e.g. BAO, SNeIa, cosmic chronometers, and so on) is inherently dependent on the assumed (late-time) cosmological model, and will in generally change if one changes the dark energy dynamics. One interesting possibility in this sense could come from measurements of the gas mass fraction $f_{\text{gas}}$ in relaxed, massive galaxy clusters, where $\approx 92$-$95\%$ of the baryons are hot and emit strongly in X-rays. This method relies only on the assumption that the matter content of rich galaxy clusters provides a fair sample of the matter content of the Universe or, in practice, that the cluster potential is able to retain all the matter within the comoving virial radius from which the cluster formed: under these assumptions, the observed ratio of baryonic to total mass should be equal to $\Omega_{b}/\Omega_{m}$, independently of any assumed cosmological model, and such a method has indeed been used over the past decades to infer $\Omega_{m}$ Allen et al. (2002); Ettori et al. (2003, 2009); Mantz et al. (2014, 2021) [^5] In practice, this method requires making implicit assumptions on General Relativity being the underlying theory of gravity when extrapolating certain scaling relations, but the dependence on the assumed model of gravity is weak. While the $f_{\text{gas}}$ method is not as widely used in the cosmological community as other probes, we believe that its minimal sensitivity to cosmological model assumptions and relatively high level of maturity compared to other less used probes should make it a very important player in the quest towards solving the Hubble tension, in the interest of calibrating $\Omega_{m}$ as model-independently as possible.
@@
In line with the previous comments, we note that our results do depend on the assumed late-time model, which here we have taken to be $\Lambda$CDM, as is usually done when studying early-time models of new physics. In line with the recent findings of Ref. Poulin et al. (2024) (see also Ref. Jedamzik et al. (2021)), we can therefore conclude that successful early-time new physics must be able to reduce $r_{d}$ while either reducing $\Omega_{m}$ or increasing $\omega_{m}$. In this sense, we agree with Ref. Poulin et al. (2024) that, when combined with early-time new physics which reduces $r_{d}$, the role of late-time new physics can be that of helping relax the constraints on either or both $\Omega_{m}$ and $S_{8}$, for instance by allowing for more freedom in the dark energy sector. We note that achieving a successful early-plus-late combination is still not a trivial feat, as not all early-time modifications follow the ideal degeneracy directions.  [^6] See for instance the explicit example in Ref. Toda et al. (2024), which fails precisely because of the varying electron mass model not following the ideal degeneracy direction for what concerns $\Omega_{m}$, as also emphasized in Ref. Poulin et al. (2024) (see also Ref. Schöneberg and Vacher (2024). Other recent case studies combining early- and late-time new physics can be found e.g. in Refs. Allali et al. (2021); Ye et al. (2023); Anchordoqui et al. (2021); Khosravi and Farhang (2022); Wang and Piao (2022); Anchordoqui et al. (2022); Reeves et al. (2023); da Costa et al. (2024); Wang and Piao (2024); Baryakhtar et al. (2024). However, as also emphasized in Ref. Poulin et al. (2024), this task is potentially made much easier by the recent DESI BAO data which, in order to be as conservative as possible, we have not adopted here.  [^7] However, we note that even when adopting DESI BAO data deviations from $\Lambda$CDM in the shape of the late-time expansion history remain constrained to $\lesssim 10\%$, as shown in the recent non-parametric analysis of Ref. Jiang et al. (2024b) by some of us. Moreover, we remark that “dark scattering” models which feature pure momentum exchange, and do not alter the background to linear order in perturbations, can be particularly promising in terms of relaxing the constraints on $\sigma_{8}$ and, potentially, $\Omega_{m}$ (see e.g. Refs. Simpson (2010); Escudero et al. (2015); Koivisto et al. (2015); Pourtsidou and Tram (2016); Kumar and Nunes (2017); Gluscevic and Boddy (2018); Boddy and Gluscevic (2018); Boddy et al. (2018); Asghari et al. (2019); Vagnozzi et al. (2020); Beltrán Jiménez et al. (2020); Becker et al. (2021); Figueruelo et al. (2021); Vagnozzi et al. (2021); Beltrán Jiménez et al. (2021); Linton et al. (2022); Rogers et al. (2022); Ferlito et al. (2022); Poulin et al. (2023); Beltrán Jiménez et al. (2023)). We believe it may be worth exploring these models in combination with successful early-time modifications, as also suggested in Ref. Vagnozzi (2023) [^8] Of course, as discussed in Ref. Vagnozzi (2023), there is the possibility that additional very-late-time or local new physics may play an important role in the Hubble tension, see e.g. Refs. Desmond et al. (2019); Ding et al. (2020); Benevento et al. (2020); Desmond and Sakstein (2020); Cai et al. (2021a); Camarena and Marra (2021); Cai et al. (2021b); Marra and Perivolaropoulos (2021); Krishnan et al. (2021b); Perivolaropoulos and Skara (2021); Castello et al. (2022); Camarena et al. (2022); Perivolaropoulos and Skara (2022b); Oikonomou et al. (2022); Högås and Mörtsell (2023a, b); Giani et al. (2024); Mazurenko et al. (2024); Huang et al. (2024); Ruchika et al. (2024b) for examples of studies in this direction. One final relevant caveat is that our entire discussion, especially the analytical arguments in Sec. II, assumed a spatially flat Universe. Relaxing this assumption is likely to relax all our results, given the well-known degeneracies between $\Omega_{m}$ and the spatial curvature parameter $\Omega_{K}$, and we plan to explore this in future work.

### v2

@@ III Datasets and methodology
We consider various combinations of the above datasets, following a rationale that will be discussed later. Note that we purposely choose not to include CMB data (not even in the form of distance priors), unlike the related Ref. Poulin et al. (2024). The reason is that we want our results to depend only on constraints arising from the late-time expansion history as much as possible, whereas including CMB data would introduce an inevitable dependence on the perturbation dynamics of the assumed early-time model, a model-dependence which we are seeking to avoid.
@@
Assuming $\Lambda$CDM as the underlying cosmological model, we explicitly investigate on real data the connection between $h$ and $\omega_{c}$ discussed earlier, see Eq. (2). We do so by first calibrating $\Omega_{m}$ by using BAO and/or SNeIa measurements, as discussed in Sec. II. Once this is achieved, we calibrate the resulting distance ladder on opposite ends, resulting in $H_{0}$ being pushed to either end of the “Hubble tension range”, $0.67\lesssim h\lesssim 0.74$. On the one hand, we can use the BBN prior on $\omega_{b}$ to infer $r_{d}$ (assuming $\Lambda$CDM), in turn calibrating BAO measurements from the early Universe side (this calibration is eventually transferred to SNeIa measurements in the same range, if they are included). The value of $H_{0}$ inferred from this inverse distance ladder is expected to fall on the low side of the Hubble tension range, i.e. closer to the Planck value Schöneberg et al. (2019, 2022b). On the other hand, the SH0ES prior on $H_{0}$ can be used to directly calibrate either or both BAO and SNeIa measurements from the local Universe side, naturally resulting in a value of $H_{0}$ falling on the high side of the Hubble tension range, i.e. closer to the SH0ES value.
@@
Finally, for purely illustrative purposes, we further quantify the relation between $\omega_{c}$, $S_{8}$, and $H_{0}$ [see Eqs. (4,14)] by fitting linear relations to the relative variations of both $\omega_{c}$ and $S_{8}$ ($\delta\omega_{c}/\omega_{c}$ and $\delta S_{8}/S_{8}$) as a function of the relative variation of $H_{0}$, $\delta h/h$:
@@ IV Results
We now present the results obtained using the methods and datasets discussed in Sec. III. Constraints on cosmological parameters of interest, as well as the coefficients $\alpha$ and $\beta$ presented in Eq. (15), are shown in Tab. 1. The following four subsections are each devoted to the comparison between the results obtained adopting two specific combinations of likelihoods discussed previously, each of which will make the mutual correlations between $\omega_{c}$, $S_{8}$, and $H_{0}$ more or less clear, while allowing us to underscore the importance of a reliable $\Omega_{m}$ calibration.
@@ IV.2 BBN+PP+BAO vs BBN+PP+SH0ES
In this case, we are no longer adopting the same dataset to calibrate $\Omega_{m}$ (earlier constrained in both cases by the BAO+PP combination), which is calibrated by BAO+PP on one side, and by PP on the other. The resulting calibration of $\Omega_{m}$ is broadly consistent, but displays some differences. In particular, from the BBN+PP+BAO combination we find $\Omega_{m}=0.314\pm 0.013$, as previously discussed in Sec. IV.1, whereas from the BBN+PP+SH0ES combination we find the $\approx 0.8\sigma$ larger value of $\Omega_{m}=0.331\pm 0.018$, as can be clearly appreciated in Fig. 2. This is not unexpected, as the fact that the PantheonPlus SNeIa sample appears to prefer slightly larger values of $\Omega_{m}\sim 0.33$ compared to its predecessors Pantheon and JLA is well known and documented in the literature Baryakhtar et al. (2024).
@@
The above increase in $H_{0}$ leads to substantial increases in both $\omega_{c}$ and $S_{8}$. In fact, we find that $\omega_{c}$ increases by $1.3\sigma$ from $0.134\pm 0.013$ to $0.159\pm 0.015$: we remark that the presence of the BBN prior in the BBN+PP+SH0ES dataset combination plays a crucial role in determining the latter value, given that PP+SH0ES on their own would only be sensitive to the sum of $\omega_{b}$ and $\omega_{c}$ (more precisely, PP is sensitive to $\Omega_{m}$, so adding the SH0ES information on $H_{0}$ naturally determines $\omega_{m}$), whereas the BBN prior allows to disentangle the baryonic contribution from the DM one. Similarly to $\omega_{c}$, we observe that $S_{8}$ increases from $0.846\pm 0.071$ to $0.980\pm 0.084$. In the case of $S_{8}$, although the central value clearly increases to extremely high values, it would be misleading to quantify the significance of the resulting tension with either CMB or weak lensing observations. In fact, we note that the uncertainty on $S_{8}$ is of order $0.07$-$0.08$, i.e. a factor of $\approx 6$-$7$ larger than that of $S_{8}=0.832\pm 0.013$ as determined by Planck, which would result in a tension formally of low significance ($\lesssim 2\sigma$) in spite of the very high central value. The reason for these large uncertainties is to be sought in the fact that our analysis includes neither CMB nor weak lensing data, both of which are crucial to reduce the uncertainty on $S_{8}$. Nevertheless, even in this way we are able to observe a significant upwards shift in $S_{8}$, in line with the analytical expectations laid out previously. Finally, the best-fit values for $\alpha$ and $\beta$ we determine for the BBN+PP+SH0ES combination are $\alpha=2.27$ and $\beta=1.73$.
@@ IV.3 BBN+BAO vs BBN+PP+SH0ES
Driven by the previous results, we now choose to compare two dataset combinations which exacerbate the tension between the respective determinations of $H_{0}$, in order to check whether a larger shift is observed in both $\omega_{c}$ and $S_{8}$. Specifically, we compare the BBN+BAO versus BBN+PP+SH0ES dataset combinations: compared to the earlier discussion in Sec. IV.2, we have removed the PP dataset from the first combination. A visual summary of our results is given in the corner plot of Fig. 3.
In this case, analogously to Sec. IV.3, we are no longer adopting the same dataset to calibrate $\Omega_{m}$, which is calibrated by the BAO dataset on one side, and the PP dataset on the other. In terms of the resulting value of $\Omega_{m}$, these two datasets display the largest difference among all the ones we discussed. In particular, from the BBN+BAO combination we find $\Omega_{m}=0.292\pm 0.019$, whereas from the BBN+PP+SH0ES combination we find the $\approx 1.5\sigma$ larger value of $\Omega_{m}=0.331\pm 0.018$ discussed previously in Sec. IV.2.
@@ IV.4 BBN+BAO vs BBN+P+SH0ES
In light of all the previous considerations, the final scenario we study is one where $\Omega_{m}$ is still calibrated by two different probes, but in a way which is more consistent than the ones considered in Sec. IV.2 and Sec. IV.3. In both cases, we saw that the use of the PantheonPlus SNeIa catalog was responsible for the higher value of $\Omega_{m}\sim 0.33$ inferred in the respective dataset combination. However, it is known that preceding Pantheon SNeIa catalog preferred lower values of $\Omega_{m}$ compared to PantheonPlus. For this reason, we now choose to compare the BBN+BAO versus BBN+P+SH0ES dataset combinations: compared to the earlier discussion in Sec. IV.3, we have therefore replaced the PP dataset in the second combination with the P one. A visual summary of our results is given in the corner plot of Fig. 4. In some sense, this comparison is the best compromise among all the ones we have considered so far. In fact, such a comparison allows for a highly consistent calibration of $\Omega_{m}$, while allowing us to explore the opposite ends of the Hubble tension range precisely because the calibration of $\Omega_{m}$ is achieved via different datasets (BAO and SNeIa), which nonetheless are consistent with each other as far as $\Omega_{m}$ is concerned.
@@ V Discussion
The results presented previously fully confirm on real data the expected correlations between $H_{0}$, $\omega_{c}$, and $S_{8}$ which, we recall, increase hand-in-hand. While we have specifically assumed $\Lambda$CDM at early times, the increase in $\omega_{c}$ and $S_{8}$ as $H_{0}$ increases does not depend on the specific model assumed for the early Universe, since it is solely a consequence of constraints on $\Omega_{m}$ imposed from the late-time expansion history  [^5] In other words, as noted in Ref. Poulin et al. (2024), uncalibrated late-time data can constrain “dimensionless” quantities such as $\Omega_{m}$, whereas a calibration is required to constrain “dimensionful” quantities such as $\omega_{c}$ – the terms dimensionless and dimensionful here are used with a slight abuse of language. In fact, repeating our analysis in a early Universe-agnostic manner, treating $r_{d}$ as a free parameter and eventually imposing some prior thereon to calibrate the BAO measurements from the early Universe side, would lead to essentially the same conclusions. Therefore, we can expect that any early-time modification to $\Lambda$CDM aiming to solve the Hubble tension will necessarily have to be accompanied by an increase in $\omega_{c}$ and, if the primordial power spectrum is $\Lambda$CDM-like, in $S_{8}$ as well. We also note that we have explicitly assumed that $\Lambda$CDM holds at late times, as is usually done when studying early-time models of new physics – we return to this point the end of the Section.
@@
A very important point which is underscored by our analysis is the key role played by a (consistent or not) $\Omega_{m}$ calibration. As highlighted especially in Sec. IV.2 and Sec. IV.3, an inconsistent calibration of $\Omega_{m}$ enhances the shifts in $\omega_{c}$ discussed in this work, which in turn exacerbates the $S_{8}$ discrepancy. This issue is actually highly relevant at present, given the mild disagreement between the values of $\Omega_{m}$ inferred from SNeIa catalogs: as we have discussed earlier, the Pantheon SNeIa dataset indicated $\Omega_{m}\sim 0.3$, which increased to $\Omega_{m}\sim 0.33$ in the PantheonPlus sample. The more recent DES-Y5 and Union3 SNeIa samples instead indicate $\Omega_{m}\sim 0.35$ Abbott et al. (2024) and $\Omega_{m}\sim 0.36$ Rubin et al. (2023) respectively, see also Refs. Baryakhtar et al. (2024); Colgáin et al. (2024a) (note that these figures are obtained within $\Lambda$CDM, and potentially increase to much larger values when allowing for an evolving dark energy component): we therefore expect that adopting these SNeIa samples would exacerbate the $\omega_{c}$ and $S_{8}$ shifts we have observed. Finally, while we defer a more detailed exploration of this point elsewhere, we note that these shifts in the values of $\Omega_{m}$ from different samples may be connected to recent discussions on redshift-evolution of inferred cosmological parameters as different redshift are probed (see e.g. Refs. Krishnan et al. (2020, 2021a); Dainotti et al. (2021, 2022a); Colgáin et al. (2022); Dainotti et al. (2022b); Colgáin et al. (2024b, 2023); Jia et al. (2023); Malekjani et al. (2024); Adil et al. (2023); Akarsu et al. (2024b); Colgáin et al. (2024c, d); Jia et al. (2024) for more details).
@@
The above discussions reinforce the urgent need, in the cosmology community, for a calibration of $\Omega_{m}$ which is as reliable and model-independent as possible. Concerning this last point, we indeed note that the value of $\Omega_{m}$ inferred from standard late-time cosmological probes (e.g. BAO, SNeIa, cosmic chronometers, and so on) is inherently dependent on the assumed (late-time) cosmological model, and will in generally change if one changes the dark energy dynamics. One interesting possibility towards a model-independent determination of $\Omega_{m}$ could come from measurements of the gas mass fraction $f_{\text{gas}}$ in relaxed, massive galaxy clusters, where $\approx 92$-$95\%$ of the baryons are hot and emit strongly in X-rays. This method relies only on the assumption that the matter content of rich galaxy clusters provides a fair sample of the matter content of the Universe or, in practice, that the cluster potential is able to retain all the matter within the comoving virial radius from which it formed: under these assumptions, the observed ratio of baryonic to total mass should be equal to $\Omega_{b}/\Omega_{m}$, independently of any assumed cosmological model, and such a method has indeed been used over the past decades to infer $\Omega_{m}$ Allen et al. (2002); Ettori et al. (2003, 2009); Mantz et al. (2014, 2021) [^6] In practice, this method requires making implicit assumptions on General Relativity being the underlying theory of gravity when extrapolating certain scaling relations, but the dependence on the assumed model of gravity is weak. While the $f_{\text{gas}}$ method is not as widely used in the cosmological community as other probes, we believe that its minimal sensitivity to cosmological model assumptions and relatively high level of maturity compared to other less used probes should make it a very important player in the quest towards solving the Hubble tension, in the interest of calibrating $\Omega_{m}$ as model-independently as possible. Another interesting possibility towards model-independently inferring $\Omega_{m}$ (with $\Omega_{b}$ known) is to use the internal properties of individual galaxies, such as stellar mass, stellar metallicity, and maximum circular velocity Villaescusa-Navarro et al. (2022); Echeverri et al. (2023); Chawak et al. (2024); Hahn et al. (2024); Wang and Pisani (2024). Nevertheless, this very interesting method, which demonstrates a potential very tight link between cosmological parameters and the astrophysics of galaxies, has yet to reach the level of maturity of all the other probes discussed so far.
@@
In line with the previous comments, we note that our results do depend on the assumed late-time model, which here we have taken to be $\Lambda$CDM, as is usually done when studying early-time models of new physics. In line with the recent findings of Ref. Poulin et al. (2024) (see also Ref. Jedamzik et al. (2021)), we can therefore conclude that successful early-time new physics must be able to reduce $r_{d}$ while either reducing $\Omega_{m}$ or increasing $\omega_{m}$ (a related result was obtained in Ref. Jedamzik et al. (2021), which however relied on a completely different argument, exploiting the different $r_{d}$-$H_{0}$ degeneracy directions for BAO versus geometrical information from the CMB). In this sense, we agree with Ref. Poulin et al. (2024) that, when combined with early-time new physics which reduces $r_{d}$, the role of late-time new physics can be that of helping relax the constraints on either or both $\Omega_{m}$ and $S_{8}$, for instance by allowing for more freedom in the dark energy sector. We note that achieving a successful early-plus-late combination is still not a trivial feat, as not all early-time modifications follow the ideal degeneracy directions.  [^7] See for instance the explicit example in Ref. Toda et al. (2024), which fails precisely because of the varying electron mass model not following the ideal degeneracy direction for what concerns $\Omega_{m}$, as also emphasized in Ref. Poulin et al. (2024) (see also Refs. Baryakhtar et al. (2024); Schöneberg and Vacher (2024)). Other recent case studies combining early- and late-time new physics can be found e.g. in Refs. Allali et al. (2021); Ye et al. (2023); Anchordoqui et al. (2021); Khosravi and Farhang (2022); Wang and Piao (2022); Anchordoqui et al. (2022); Reeves et al. (2023); da Costa et al. (2024); Wang and Piao (2024); Baryakhtar et al. (2024). However, as also emphasized in Ref. Poulin et al. (2024) (see also Ref. Baryakhtar et al. (2024)), this task is potentially made much easier by the recent DESI BAO data which, in order to be as conservative as possible, we have not adopted here.  [^8] However, we note that even when adopting DESI BAO data deviations from $\Lambda$CDM in the shape of the late-time expansion history remain constrained to $\lesssim 10\%$, as shown in the recent non-parametric analysis of Ref. Jiang et al. (2024b) by some of us. Moreover, we remark that “dark scattering” models which feature pure momentum exchange, and do not alter the background to linear order in perturbations, can be particularly promising in terms of relaxing the constraints on $\sigma_{8}$ and, potentially, $\Omega_{m}$ (see e.g. Refs. Simpson (2010); Escudero et al. (2015); Koivisto et al. (2015); Pourtsidou and Tram (2016); Kumar and Nunes (2017); Gluscevic and Boddy (2018); Boddy and Gluscevic (2018); Boddy et al. (2018); Asghari et al. (2019); Vagnozzi et al. (2020); Beltrán Jiménez et al. (2020); Becker et al. (2021); Figueruelo et al. (2021); Vagnozzi et al. (2021); Beltrán Jiménez et al. (2021); Linton et al. (2022); Rogers et al. (2022); Ferlito et al. (2022); Poulin et al. (2023); Beltrán Jiménez et al. (2023)). We believe it may be worth exploring these models in combination with successful early-time modifications, as also suggested in Ref. Vagnozzi (2023) [^9] Of course, as discussed in Ref. Vagnozzi (2023), there is the possibility that additional very-late-time or local new physics may play an important role in the Hubble tension, see e.g. Refs. Desmond et al. (2019); Ding et al. (2020); Benevento et al. (2020); Desmond and Sakstein (2020); Cai et al. (2021a); Camarena and Marra (2021); Cai et al. (2021b); Marra and Perivolaropoulos (2021); Krishnan et al. (2021b); Perivolaropoulos and Skara (2021); Castello et al. (2022); Camarena et al. (2022); Perivolaropoulos and Skara (2022b); Oikonomou et al. (2022); Högås and Mörtsell (2023a, b); Giani et al. (2024); Mazurenko et al. (2024); Huang et al. (2024); Ruchika et al. (2024b) for examples of studies in this direction. One final crucial caveat is that our entire discussion, especially the analytical arguments in Sec. II, assumed a spatially flat Universe. Relaxing this assumption is likely to relax all our results, given the well-known degeneracies between $\Omega_{m}$ and the spatial curvature parameter $\Omega_{K}$, and we plan to explore this in future work.